freshavacado850:

pls help I have 3 questions

freshavacado850:

hold up lemme get a ss

freshavacado850:

1 attachment
freshavacado850:

1 attachment
freshavacado850:

1 attachment
Lura:

1:Answer A 2:B 3:C

TETSXPREME:

The second question 3 x -1 =8 Your supposed to add 1 to each of the sides 3 𝑥−1+1=8+1

TETSXPREME:

3𝑥−1+1=8+1 *

freshavacado850:

oh

TETSXPREME:

Do you understand it ?

freshavacado850:

no

freshavacado850:

that was the third one

TETSXPREME:

You see the end of each equation

freshavacado850:

yes